Through how many volts of potential difference must an electron be accelerated to achieve a wavelength of 0.27 nm?
Answer #1

if V (volts) is the potential diff by which "e" was accelerated, gain in its kinetic energy

KE=\frac{mv^{2}}{2}=eV
v=\sqrt{\frac{2eV}{m}}

mv=\sqrt{2eVm}
  
wavelength associated with electron

\lambda =\frac{h}{mv}

\lambda =\frac{h}{\sqrt{2eVm}}
\sqrt{2eVm}=\frac{h}{\lambda }

V=\frac{(\frac{h}{\lambda })^{2}}{2em}
V=\frac{(\frac{6.6\cdot 10^{-34}}{0.27\cdot 10^{-9} })^{2}}{2\cdot 9.1\cdot 10^{-31}\cdot 1.6\cdot 10^{-19}}

V=20.51\, V